Enter your answer in the provided box The osmotic pressure of 1.74 × 10-2 M solutions of CaCl2 and urea at 25°C are 1.05 and 0.425 atm, respectively. Calculate the vant Hoff factor for the CaCl solution References eBook & Resources
Answer #1

Concentration of CaCl2 solution, C-1.74×10-2 M = 1.74×10-2 mol/L Gas constant, S = 0.0821 L atm /K·mol Osmotic pressure, 1.05